# Exploiting Unlabeled Data with Multiple Expert Teachers for Open Vocabulary Aerial Object Detection and Its Orientation Adaptation

### v1

@@ V-B Ablation studies
Background Embeddings
We compare three common background embeddings: fixed all-zero embedding[35], fixed mean embedding[96], and learnable background embedding[97]. As shown in Table VII, the choice of background embedding has slight impact on our experimental results, with the learnable embedding showing better performance.
@@
We evaluate four types of prompt templates, including T1: “[category]”, T2: “a [category]”, T3: “a satellite photo of [category]”, and T4: “a photo of [category]”. According to Table VIII, T4 is more suitable.
Function $\mathrm{f}(\cdot)$ of AJV Score
Table IX compares two types of functions $\mathrm{f}(\cdot)$ in Equ.(16), i.e., $\mathrm{Identity}(\cdot)$ and $\sin(\cdot)$. The results indicate that using $\sin(\cdot)$ function for angle transformation yields better performance. This improvement is attributed to the fact that the oriented box representation exhibits discontinuities at boundaries, and the $\sin(\cdot)$ function helps to smooth these discontinuities, resulting in more reliable variance estimates.
TABLE VIII: Ablation with different prompt templates.
@@
TABLE IX: Ablation on Function $\mathrm{f}(\cdot)$ of AJV Score.
@@
TABLE X: Boost GLIP/GroundingDINO with our pseudo-labeling strategies. $\ddagger$: Use pseudo-labels generated by our pseudo-labeling strategies to finetune the model.
Method mAP mAPbase mAPnovel HM
GLIP 57.5 65.2 8.0 14.2
GroundingDINO 58.0 66.0 5.8 10.6
GLIP 59.7 66.6 14.6 24.0
GroundingDINO 47.0 51.5 17.6 26.3
TABLE XI: Evaluation on CLIPs (e.g., RemoteCLIP, CLIP) with ground-truth (GT) or RPN proposals.
@@
TABLE XII: Horizontal open-vocabulary detection benchmark on DIOR dataset. $\mathcal{T}_{\mathrm{novel}}$ indicates whether the novel classes need to be pre-known. $\mathcal{I}_{\mathrm{cls}}$ denotes whether extra classification or caption datasets are required during training. $\dagger$: The best result on the zero-shot object detection leaderboard of VisDrone2023. $\ddagger$: The results of our own implementation. The short names for categories are defined as: Airport (APO), Basketball court (BC), Ground track field (GTF) and Windmill (WM).
@@
ViLD [34] Faster R-CNN R50 9.1 11.8 26.9 9.1 25.6 28.5 14.2 19.0
OV-DETR [85] Deformable DETR R50 29.9 20.5 26.8 5.2 28.7 30.8 20.6 24.7
Detic [24] CenterNet2 R50 2.0 2.2 15.0 0.0 16.8 19.8 4.8 7.7
@@
GLIP [19] GLIP Swin-T 0.0 3.3 9.1 9.1 33.8 41.0 5.4 9.5
BARON [98] Faster R-CNN R50 22.8 18.3 34.0 3.0 27.4 29.4 19.5 23.5
YOLO-World [99] YOLOv8-M YOLOv8-M 0.0 20.1 13.7 0.0 32.9 39.1 8.5 13.9
@@
ViLD [34] Faster R-CNN R50 9.1 11.8 26.7 9.1 45.7 53.5 14.2 22.4
OV-DETR [85] Deformable DETR R50 29.1 19.7 25.6 5.2 54.0 62.6 19.9 30.2
Detic [24] CenterNet2 R50 0.7 2.2 11.2 0.0 36.9 45.3 3.5 6.5
GroundingDINO [21] GroundingDINO R50 0.3 3.3 9.1 0.2 57.3 70.8 3.2 6.2
GLIP [19] GLIP Swin-T 0.0 2.4 9.1 3.0 56.0 69.1 3.6 6.9
BARON [98] Faster R-CNN R50 14.1 16.6 30.4 0.1 50.9 59.8 15.3 24.4
@@
TABLE XIII: Oriented open-vocabulary detection benchmark on DOTA-v1.0 dataset. The short names for categories are defined as: Plane (PL), Baseball diamond (BD), Ground track field (GTF) and Basketball court (BC).
@@
ViLD [34] Faster-RCNN R50 88.3 64.7 5.5 0.4 55.1 63.2 2.9 5.6
@@
GLIP [19] GLIP R50 86.8 64.6 9.8 6.1 57.5 65.2 8.0 14.2
GLIP [19] GLIP Swin-T 71.9 20.5 3.1 5.6 31.4 35.6 4.3 7.7
GroundingDINO [21] GroundingDINO R50 88.4 53.2 3.0 0.7 53.1 61.0 1.9 3.6
@@
TABLE XIV: Oriented open-vocabulary detection benchmark on DIOR dataset. The short names for categories are defined as: Airplane (APL), Airport (APO), Basketball court (BC), Ground track field (GTF) and Windmill (WM).
@@
ViLD [34] Oriented-RCNN R50 53.6 0.2 2.3 16.3 9.1 39.3 47.4 7.0 12.2
GLIP [19] GLIP R50 43.3 0.1 1.6 20.2 0.0 39.4 47.9 5.5 9.8
GLIP [19] GLIP Swin-T 33.3 0.1 2.6 13.8 0.0 28.7 34.8 4.1 7.4
GroundingDINO [21] GroundingDINO R50 43.7 0.1 0.4 10.8 0.0 39.5 48.7 2.8 5.3
@@
TABLE XV: Results on open-vocabulary COCO. $\dagger$: Results quoted from [81]. $\ddagger$: The results of our own implementation, under the same experimental setup as ours. *: Results quoted from the original paper.
mAPnovel mAPbase mAP50all HM
Mask R-CNN [40] - 51.4 - -
WSDDN [100$\dagger$ 19.7 19.6 19.6 19.6
Cap2Det [101$\dagger$ 20.3 20.1 20.1 20.2
OVR-CNN [81] 22.8 46.0 39.9 30.5
Detic [24] 24.1 52.0 44.7 32.9
ViLD [34$\ddagger$ 12.9 38.8 32.0 19.3
PromptDet [84] 26.6 59.1 50.6 36.7
CastDet (Ours) 30.3 47.4 42.9 37.0
Cross evaluation.
Some popular detectors for natural images, such as GroundingDINO and GLIP, underperform on novel aerial categories due to a lack of a robust strategy for discovering novel aerial objects. Thus, we provide them with pseudo-labels filtered by our box selection strategies, i.e., AJV and RJV. As shown in Table X, this leads to significant improvements, e.g., 6.6% mAPnovel gains for GLIP and 11.8% mAPnovel gains for GroundingDINO, highlighting the effectiveness of our approach in bridging the gap between general object detection frameworks and the specialized requirements of aerial imagery.
V-C Comparisons with state-of-the-art methods
V-C1 Horizontal Object Detection
Comparison with CLIPs
We compare our method with RemoteCLIP [33] under two settings: (1) using ground truth bounding boxes as proposals to measure open-vocabulary classification performance, and (2) employing proposals predicted by RPN to assess overall model performance. Table XI shows that CastDet outperforms RemoteCLIP in open-vocabulary classification, achieving an improvement from 37.6% to 47.3% mAP. Furthermore, CastDet shows significant performance gains when using RPN proposals, with mAP increasing from 11.6% to 38.1%. Additional experiments conducted on COCO [16] further validate the improvements achieved by our approach on natural images.
Comparison with SOTAs
In OVD, auxiliary weak supervision may be employed in various forms: 1) leveraging both base and novel categories ($\mathcal{T}_{\mathrm{novel}}$) with unlabeled images for knowledge distillation or pseudo-labeling, such as ViLD [34] and OV-DETR [85], which is consistent with our setting; and 2) using image classification or image caption datasets ($\mathcal{I}_{\mathrm{cls}}$) that contain novel concepts, such as Detic [24], etc [99, 21, 19, 98]. To support these approaches, we provide the NWPU-RESISC45 [93] classification dataset. We conduct experiments on DIOR and VisDroneZSD, as shown in Table XII. Our method outperforms the previous SOTA by 26.8% mAPnovel. Visualization results are presented in Fig. 11.
V-C2 Oriented Object Detection
Given the absence of an oriented open-vocabulary detector, we adhere to the architectures used for horizontal detection and extend them for oriented detection. We compare the proposed method with Oriented ViLD, Oriented GLIP and Oriented GroundingDINO. A benchmark for oriented open-vocabulary detection is established on two oriented datasets: DOTA-v1.0 and DIOR, with results shown in Table XIII and Table XIV. Our method achieves competitive performance, e.g., 36.0 % mAPnovel on DOTA-v1.0 and 24.3% mAPnovel on DIOR. Fig. 11 shows the visualization results.
V-C3 Natural Images Open-vocabulary Detection
To further assess the effectiveness of the proposed method beyond aerial imagery, we extend our evaluations to include the COCO dataset. As shown in Table XV, our method achieves impressive results in novel category detection, reaching a 30.3% mAPnovel. The qualitative results are shown in Fig. 12.
VI Conclusion
@@
In this paper, we address open-vocabulary aerial object detection, tackling challenges such as limited annotated data and unique characteristics in aerial scenarios. We present CastDet, a CLIP-activated student-teacher detector that supports horizontal and oriented detection. By incorporating a flexible student-(multi)teacher self-learning approach, a dynamic label queue, and several box selection strategies tailored for aerial scenarios, our method effectively bridges the gap between general open-vocabulary detection and the specific needs of aerial imagery. We conduct experiments on several aerial datasets, such as DIOR and DOTA, achieving significant performance improvements. To the best of our knowledge, this marks the first effort in open-vocabulary aerial detection. We aspire for the proposed techniques and the established benchmark to lay a solid foundation for future research in this area.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (62071333), Fundamental Research Funds for the Central Universities (22120220654), Tongji University national artificial intelligence production-education integration innovation project, and the Science and Technology on Near-Surface Detection Laboratory (6142414221606).

### v2

@@ 4.2 Ablation studies
We evaluate four types of prompt templates, including T1: “[category]”, T2: “a [category]”, T3: “a satellite photo of [category]”, and T4: “a photo of [category]”. According to Table 10, T4 is more suitable.
Table 10: Ablation with different prompt templates on DOTA.
@@
Function $\mathrm{f}(\cdot)$ of AJV Score
Table 11 compares two types of functions $\mathrm{f}(\cdot)$ in Equ.(16), i.e., $\mathrm{Identity}(\cdot)$ and $\sin(\cdot)$. The results indicate that using $\sin(\cdot)$ function for angle transformation yields better performance. This improvement is attributed to the fact that the oriented box representation exhibits discontinuities at boundaries, and the $\sin(\cdot)$ function helps to smooth these discontinuities, resulting in more reliable variance estimates.
Table 11: Ablation on Function $\mathrm{f}(\cdot)$ of AJV Score on DOTA.
@@
Computation Costs.
Table 12 compares the computational costs of various methods in terms of FLOPs, parameter count, memory consumption, and inference speed (FPS), measured on an NVIDIA A6000 GPU. During training, CastDet utilizes teacher models to guide the student model training. After training, both the teacher and the student models are capable of detecting objects independently. Therefore, CastDet demonstrates superior efficiency, with the lowest parameter count (42.17M) and the highest FPS (9.8).
Table 12: Comparison of computational costs across different methods.
[…]
GLIP 89.66 147.70 568 5.4
GroundingDINO 243.22 172.15 662 6.2
[…]
4.3 Comparisons with state-of-the-art methods
Table 13: Evaluation on CLIPs (e.g., RemoteCLIP, CLIP) with ground-truth (GT) or RPN proposals.
@@ 4.3 Comparisons with state-of-the-art methods
Table 14: Horizontal open-vocabulary detection benchmark on DIOR dataset. $\mathcal{T}_{\mathrm{novel}}$ indicates whether the novel classes need to be pre-known. $\mathcal{I}_{\mathrm{cls}}$ denotes whether extra classification or caption datasets are required during training. $\dagger$: The best result on the zero-shot object detection leaderboard of VisDrone2023. $\ddagger$: The results of our own implementation. The short names for categories are defined as: Airport (APO), Basketball court (BC), Ground track field (GTF) and Windmill (WM). S and LT denote the student model and the localization teacher model respectively.
@@
ViLD [gu2021open] Faster R-CNN R50 9.1 11.8 26.9 9.1 25.6 28.5 14.2 19.0
OV-DETR [zang2022open] Deformable DETR R50 29.9 20.5 26.8 5.2 28.7 30.8 20.6 24.7
Detic [zhou2022detic] CenterNet2 R50 2.0 2.2 15.0 0.0 16.8 19.8 4.8 7.7
@@
GLIP [9879567glip] GLIP R50 0.1 5.7 14.9 0.0 28.3 34.9 5.2 9.0
GLIP [9879567glip] GLIP Swin-T 0.0 3.3 9.1 9.1 33.8 41.0 5.4 9.5
BARON [wu2023baron] Faster R-CNN R50 22.8 18.3 34.0 3.0 27.4 29.4 19.5 23.5
YOLO-World [Cheng2024YOLOWorld] YOLOv8-M YOLOv8-M 0.0 20.1 13.7 0.0 32.9 39.1 8.5 13.9
@@
ViLD [gu2021open] Faster R-CNN R50 9.1 11.8 26.7 9.1 45.7 53.5 14.2 22.4
OV-DETR [zang2022open] Deformable DETR R50 29.1 19.7 25.6 5.2 54.0 62.6 19.9 30.2
Detic [zhou2022detic] CenterNet2 R50 0.7 2.2 11.2 0.0 36.9 45.3 3.5 6.5
GroundingDINO [liu2023groundingdino] GroundingDINO R50 0.3 3.3 9.1 0.2 57.3 70.8 3.2 6.2
GLIP [9879567glip] GLIP R50 0.0 2.1 13.4 0.0 50.3 63.6 3.9 7.3
GLIP [9879567glip] GLIP Swin-T 0.0 2.4 9.1 3.0 56.0 69.1 3.6 6.9
BARON [wu2023baron] Faster R-CNN R50 14.1 16.6 30.4 0.1 50.9 59.8 15.3 24.4
@@
Table 15: Oriented open-vocabulary detection benchmark on DOTA-v1.0 dataset. The short names for categories are defined as: Plane (PL), Baseball diamond (BD), Ground track field (GTF) and Basketball court (BC).
@@
ViLD [gu2021open] Faster-RCNN R50 88.3 64.7 5.5 0.4 55.1 63.2 2.9 5.6
@@
GLIP [9879567glip] GLIP R50 86.8 64.6 9.8 6.1 57.5 65.2 8.0 14.2
GLIP [9879567glip] GLIP Swin-T 71.9 20.5 3.1 5.6 31.4 35.6 4.3 7.7
GroundingDINO [liu2023groundingdino] GroundingDINO R50 88.4 53.2 3.0 0.7 53.1 61.0 1.9 3.6
@@
Table 16: Oriented open-vocabulary detection benchmark on DIOR dataset. The short names for categories are defined as: Airplane (APL), Airport (APO), Basketball court (BC), Ground track field (GTF) and Windmill (WM).
@@
ViLD [gu2021open] Oriented-RCNN R50 53.6 0.2 2.3 16.3 9.1 39.3 47.4 7.0 12.2
GLIP [9879567glip] GLIP R50 43.3 0.1 1.6 20.2 0.0 39.4 47.9 5.5 9.8
GLIP [9879567glip] GLIP Swin-T 33.3 0.1 2.6 13.8 0.0 28.7 34.8 4.1 7.4
GroundingDINO [liu2023groundingdino] GroundingDINO R50 43.7 0.1 0.4 10.8 0.0 39.5 48.7 2.8 5.3
@@
Table 17: Oriented open-vocabulary detection benchmark on STAR dataset. The short names for categories are defined as: Ship (SH), Airplane (PL), Cooling tower (CT), and Basketball court (BC).
Method Detector Backbone $\mathcal{T}_{\mathrm{novel}}$ $\mathcal{I}_{\mathrm{cls}}$ SH PL CT BC mAP mAPbase mAPnovel HM
ViLD [gu2021open] Faster-RCNN R50 37.1 80.6 30.6 4.5 16.1 18.4 4.4 7.1
[…]
GLIP [9879567glip] GLIP R50 27.1 79.3 9.7 7.5 13.4 15.1 4.8 7.3
GLIP [9879567glip] GLIP Swin-T 0.5 21.2 0.4 0.6 1.4 1.6 0.2 0.4
GroundingDINO [liu2023groundingdino] GroundingDINO R50 37.3 79.6 16.3 11.9 18.6 21.3 5.3 8.5
GroundingDINO [liu2023groundingdino] GroundingDINO Swin-T 38.1 79.2 33.5 12.5 17.8 19.9 7.6 11.0
CastDet (S) Faster-RCNN R50 37.3 80.3 25.6 18.8 16.4 18.4 6.4 9.5
[…]
4.3.1 Horizontal Object Detection
Comparison with CLIPs
We compare our method with RemoteCLIP [liu2023remoteclip] under two settings: i) using ground truth bounding boxes as proposals to measure open-vocabulary classification performance, and ii) employing proposals predicted by RPN to assess overall model performance. Table 13 shows that CastDet outperforms RemoteCLIP in open-vocabulary classification, achieving an improvement from 37.6% to 47.3% mAP. Furthermore, CastDet shows significant performance gains when using RPN proposals, with mAP increasing from 11.6% to 38.1%. Additional experiments conducted on COCO [lin2014microsoft] further validate the improvements achieved by our approach on natural images.
Comparison with SOTAs
In OVD, auxiliary weak supervision may be employed in various forms: 1) leveraging both base and novel categories ($\mathcal{T}_{\mathrm{novel}}$) with unlabeled images for knowledge distillation or pseudo-labeling, such as ViLD [gu2021open] and OV-DETR [zang2022open], which is consistent with our setting; and 2) using image classification or image caption datasets ($\mathcal{I}_{\mathrm{cls}}$) that contain novel concepts, such as Detic [zhou2022detic], etc [Cheng2024YOLOWorld, liu2023groundingdino, 9879567glip, wu2023baron]. To support these approaches, we provide the NWPU-RESISC45 [NWPU_RESISC45_7891544] classification dataset. We conduct experiments on DIOR and VisDroneZSD, as shown in Table 14. Our method outperforms the previous SOTA by 26.8% mAPnovel. Visualization results are presented in Fig. 13.
Comparison with Semi-supervised Methods
Table 18 compares CastDet with representative semi-supervised object detection methods, including Instant-Teaching [zhou2021instant], Unbiased Teacher [liu2021unbiased], and Soft-Teacher [softteacher_xu2021end]. While these methods achieve reasonable recall on base categories, their performance on novel categories completely collapses (mAPnovel). In contrast, CastDet improves both recall and precision, achieving 40.3 mAPnovel and 39.1 HM, while also boosting ARnovel to 69.1. These results highlight the clear advantage of CastDet for open-vocabulary detection, especially in discovering novel categories.
Table 18: Comparison with semi-supervised methods on VisDroneZSD.
Method ARall ARbase ARnovel HM mAP mAPbase mAPnovel HM
Instant-Teaching [zhou2021instant] 60.0 60.0 61.6 60.5 23.2 29.0 0.0 0.0
Unbiased Teacher [liu2021unbiased] 64.2 64.4 63.7 64.0 24.0 30.1 0.0 0.0
Soft-Teacher [softteacher_xu2021end] 60.1 63.2 47.7 54.4 28.7 35.9 0.0 0.0
CastDet(Ours) 63.6 62.2 69.1 65.5 38.4 37.9 40.3 39.1
4.3.2 Oriented Object Detection
Given the absence of an oriented open-vocabulary detector, we adhere to the architectures used for horizontal detection and extend them for oriented detection. We compare the proposed method with Oriented ViLD, Oriented GLIP and Oriented GroundingDINO. A benchmark for oriented open-vocabulary detection is established on three oriented datasets: DOTA-v1.0, DIOR, and STAR, with results shown in Table 15, Table 16 and Table 17. Our method achieves competitive performance, e.g., 36.0 % mAPnovel on DOTA-v1.0, 24.3% mAPnovel on DIOR and 19.4% mAPnovel on STAR. Fig. 13 shows the visualization results.
4.3.3 Natural Images Open-vocabulary Detection
To further assess the effectiveness of the proposed method beyond aerial imagery, we extend our evaluations to include the COCO dataset. As shown in Table 19, our method achieves impressive results in novel category detection, reaching a 30.3% mAPnovel. The qualitative results are shown in Fig. 14.
Table 19: Results on open-vocabulary COCO. $\dagger$: Results quoted from [zareian2021open]. $\ddagger$: The results of our own implementation, under the same experimental setup as ours. *: Results quoted from the original paper.
mAPnovel mAPbase mAP50all HM
Mask R-CNN [maskrcnn_he2017mask] - 51.4 - -
[…]
OVR-CNN [zareian2021open] 22.8 46.0 39.9 30.5
Detic [zhou2022detic] 24.1 52.0 44.7 32.9
[…]
PromptDet [feng2022promptdet] 26.6 59.1 50.6 36.7
CastDet (Ours) 30.3 47.4 42.9 37.0
@@ 4.3.3 Natural Images Open-vocabulary Detection
5 Discussion and Conclusion
@@
In this paper, we address open-vocabulary aerial object detection, tackling challenges such as limited annotated data and unique characteristics in aerial scenarios. We present CastDet, a CLIP-activated student-teacher detector that supports both horizontal and oriented detection. By incorporating a flexible student-(multi)teacher self-learning approach, a dynamic label queue, and several box selection strategies, our method effectively bridges the gap between general open-vocabulary detection and the specific needs of aerial imagery.
Beyond its strong performance on two-stage detectors, such as Faster R-CNN and Oriented R-CNN, CastDet also exhibits potential for adaptation to other detection frameworks. Specifically, by adjusting the representation of the confidence score, our method can be transferred to one-stage detectors, where the final confidence score can serve as a substitute for the RPN score. Likewise, for DETR-based detectors, scores such as RJV, BJV, SJV, and AJV can be derived by computing the variance of regression predictions across different transformer layers, providing an effective measure of prediction uncertainty.
Extensive experiments on multiple aerial datasets, including VisDroneZSD, DIOR and DOTA, validate that CastDet achieves significant performance improvements. To the best of our knowledge, this marks the first effort in open-vocabulary aerial detection. We aspire for the proposed techniques and the established benchmark to lay a solid foundation for future research in this area.